# Neural Empirical Interpolation Method for Nonlinear Model Reduction

### v1

@@ 3.4.1. Local Error Estimate
We now seek a theoretical estimate to explain the sources of error in NEIM. For simplicity, we will consider the case in which the NEIM networks approximate constant vectors, that is, $w_{t}^{(j)}(\mu_{i})>0$ if and only if $\mu_{i}=\mu^{(j)}$. We will also assume for simplicity that $w_{e}(\mu_{i};\mu_{j})=\delta_{ij}$.
Define the matrix of exact vectors
$M^{(k)}=\begin{bmatrix}\vline&&\vline\\ z^{(1)}(\mu^{(1)})&\dots&z^{(k)}(\mu^{(k)})\\ \vline&&\vline\end{bmatrix},$
and the matrix of its approximation by neural networks,
$\widehat{M}^{(k)}(\mu)=\begin{bmatrix}\vline&&\vline\\ M_{\mu^{(1)}}(\tilde{v}(\mu))&\dots&M_{\mu^{(k)}}(\tilde{v}(\mu))\\ \vline&&\vline\end{bmatrix}.$
Now define the vector of interpolated $\theta$ coefficients resulting from NEIM to be
@@
With a litte abuse of notation we denote $\operatorname{N}(\mu):=U^{\top}\operatorname{N}(v(\mu);\mu)$, and we wish to estimate
@@
$\leq\|\operatorname{N}-M^{(k)}\theta^{(k)}\|_{L^{2}}+\sum_{i=1}^{k}\|z^{(i)}(\mu^{(i)})-M_{\mu^{(i)}}(\tilde{v}(\cdot))\|_{L^{2}}\|\theta_{i}^{(k)}\|_{L^{2}}+\sum_{i=1}^{k}\|M_{\mu^{(i)}}(\tilde{v}(\cdot))\|_{L^{2}}\|\theta_{i}^{(k)}-\widehat{\theta}_{i}^{(k)}\|_{L^{2}}$
$\leq\underbrace{\|\operatorname{N}-M^{(k)}\theta^{(k)}\|_{L^{2}}}_{\text{Projection error}}+\sum_{i=1}^{k}\underbrace{\|z^{(i)}(\mu^{(i)})-M_{\mu^{(i)}}(\tilde{v}(\cdot))\|_{L^{2}}}_{\text{Training error}}\|\theta_{i}^{(k)}\|_{L^{2}}+\lvert\mathcal{P}\rvert^{1/2}\sum_{i=1}^{k}\underbrace{\|\theta_{i}^{(k)}-\widehat{\theta}_{i}^{(k)}\|_{L^{2}}}_{\text{Interpolation error}}$
$\hskip 21.52771pt+\sum_{i=1}^{k}\underbrace{\|M_{\mu^{(i)}}(\tilde{v}(\cdot))-z^{(i)}(\mu^{(i)})\|_{L^{2}}\|\theta_{i}^{(k)}-\widehat{\theta}_{i}^{(k)}\|_{L^{2}}}_{\text{Training/interpolation error interaction}},$
where in the last inequality, we used the triangle inequality and the fact that $\|z^{(i)}(\mu^{(i)})\|_{2}=1$.
@@ 3.5. Comparison with State-of-the-Art Techniques
In this section, we qualitatively compare NEIM and the discrete empirical interpolation method (DEIM) [8] to highlight a few advantages of NEIM. We then briefly relate NEIM to DeepONets [26].
@@ 4. Numerical examples
We now present some numerical examples to demonstrate the performance of NEIM. Code for all examples is available at https://github.com/maxhirsch/NEIM. For each experiment, we use PyTorch’s learning rate scheduler and the Adam optimizer for training the NEIM networks. The other parameters for the training and NEIM network architecture for the examples are listed in Table 2.
@@
For the liquid crystal experiment, our two NEIM approximations used the same training parameters, which we list in the “Liquid Crystal $Q$, $r$” row. The “Liquid Crystal $\theta$” row in the table lists the parameters for training the neural network which outputs the prediction of $\theta(\mu)$. The other experiments use linear interpolation for the prediction of $\theta(\mu)$.
We use the open source library MLniCS [27], based on RBniCS [36] and FEniCS [25], for our physics-informed neural network example.
In what follows, the “number of modes” used in DEIM will refer to the number of components of the nonlinearity which DEIM selects to interpolate. In NEIM, the “number of modes” will simply refer to the number of terms in the NEIM expansion.
@@ 4.1. Solution Independent Function Approximation
In this first example, we approximate a parameterized constant function that can be thought of as the forcing term in a parameterized PDE discretization. Let us consider $s:\Omega\times\mathcal{P}\to\mathbb{R}$ defined by
$s(x;\mu)=(1-x)\cos(3\pi\mu(x+1))e^{-(1+x)\mu},$
where $\Omega=[-1,1]$ and $\mu\in\mathcal{P}=[1,\pi]$. Then let $\boldsymbol{x}=[x_{1},\dots,x_{n}]^{\top}\in\mathbb{R}^{n}$ with $x_{i}$ equally spaced points in $\Omega$ for $i=1,\dots,n$, with $n=100$, and define the forcing $f:\mathcal{P}\to\mathbb{R}^{n}$ by
$f(\mu)=[s(x_{1};\mu),\dots,s(x_{n};\mu)]^{\top}\in\mathbb{R}^{n}$
for $\mu\in\mathcal{P}$. The goal is to solve the equation
@@
coming from a finite difference approximation of a second order problem where $v_{1}=v_{n}=0$, $h^{-2}=30$, and
@@
We now compute NEIM approximation, where we take all weights equal to $1$, and compare with DEIM approach. Because the NEIM neural networks in this case approximate a vector at each iteration, we also include a comparison to “exact” NEIM, that is, the result of NEIM when the neural networks are replaced with the exact constant vectors. This way, we can analyze the performance of NEIM in relation to the accuracy of the different components in the expansion.
In Figure 3, we see how the error in NEIM and DEIM develop as the number of modes used increases. The number of modes in DEIM is the number of singular vectors retained in the DEIM POD of the nonlinearity, while the number of modes in NEIM is the number of neural networks in the NEIM expansion. The error for all methods decreases with respect to the number of modes used. For larger numbers of modes, reaching machine precision for the singular values, DEIM outperforms NEIM, but when considering only fewer modes, NEIM performance are remarkably great over the whole parameter space. Indeed, for $r<20$ the neural empirical interpolation method shows small error at the parameters that have been selected, outperforming DEIM for many parameter values. DEIM seems to have roughly the same error across the parameter space for each number of modes used, while NEIM shows error of different order of magnitude. As expected, the error for the exact version of NEIM, exploiting the exact constant vectors as modes, is smaller than that of the normal NEIM when using a large number of modes. Despite this, we remark that the accuracy of NEIM and NEIM “exact” are equivalent for $r<20$, confirming the consistency of the approach.
@@
Figure 3. NEIM and DEIM absolute error by parameter for increasing number of modes in experiment 4.1.
@@
Figure 4. NEIM and DEIM average absolute error by number of modes in experiment 4.1.
@@ 4.2. Solution Dependent Function Approximation
We now approximate the parameterized nonlinear function $s:\Omega\times\mathbb{R}\times\mathcal{P}\to\mathbb{R}$ defined by
$s(x,v;\mu)=(1-\lvert x\rvert)e^{-(1+x)v\mu},$
where $\Omega=[-1,1]$ and $\mu\in\mathcal{P}=[1,\pi]$, and $x_{i}$ are equally spaced points in $\Omega$ for $i=1,\dots,n$, with $n=100$. Let $v(\mu)\in\mathbb{R}^{n}$ be the solution of the system,
@@
where $h^{-2}=30$ and $f(v;\mu)$ is the nonlinear term defined by
$f(v;\mu)=\begin{bmatrix}s(x_{1},v_{1};\mu)&\dots&s(x_{n},v_{n};\mu)\end{bmatrix}^{\top}.$
As before, we use $m=51$ equally spaced parameters in $\mathcal{P}=[1,\pi]$, and perform a POD on the matrix of solution snapshots to obtain a POD basis $U_{r}$. The goal now is to approximate $U_{r}^{\top}f(U_{r}\tilde{v};\mu)$ for $\mu\in\mathcal{P}$. Based on the plot of singular values in Figure 5, we truncate our expansion at $r=20$ modes.
@@
For this test case, we take the weights in the NEIM algorithm to be
$w_{e}(\mu_{i};\mu_{j})=\delta_{ij},\quad w_{t}^{(k)}(\mu_{i})=1\quad\forall i,j,k,$
and compare the accuracy results of NEIM and DEIM approaches.
@@
Figure 7. NEIM and DEIM average absolute error by number of modes in experiment 4.2.
@@ 4.3. Nonlinear Elliptic Physics-Informed Neural Network
We now apply NEIM in the context of physics-informed neural networks (PINNs) for reduced order models. The goal is to combine a physics-based learning within a data-driven reduced context. More specifically, one aims at learning the parameter-to-reduced vector mapping via neural network architectures that minimize a joint loss given by a supervised term, accounting for projection onto a linear POD subspace, and an unsupervised term, based on the reduced residual projected onto the same base. For further details we remind the reader to [20, 9, 30]. In order to embed the reduced physics in the loss for nonlinear problems, it is important that the approximation of the nonlinearity is compatible with automatic differentiation. When it comes to more complex nonlinearities, exploiting tensor-assembly for polynomial nonlinearities or empirical interpolation approaches can be challenging. In fact, DEIM will require to assemble nonlinearities in the online phase of computation, which, if done in standard finite element packages like FEniCS [25], will not be differentiable in deep learning libraries like PyTorch [29]. On the contrary, NEIM is automatically compatible with deep learning libraries because our approach is itself a neural network with automatic differentiation included by definition.
@@
We exploit the finite element method as high fidelity solver to discretize the PDE. Then, we perform a POD on $m=100$ equispaced points in parameter space to obtain $U_{r}$. In this case, our goal is to approximate $U_{r}^{\top}f(U_{r}\tilde{v};\mu)$, where $f$ is given by
$f(v;\mu)_{i}=\frac{\mu_{1}}{\mu_{2}}\left(\exp\left(\mu_{2}v_{i}\right)-1\right)\Delta x_{i},\quad i=1,\dots,n,$
and $\Delta x_{i}$ is the volume of the finite element basis test function at node $i$. In other words, the nonlinearity we aim at approximating is a low order quadrature rule for the nonlinear integral term in the FE formulation of the problem. The weights we use for NEIM are given by $w_{e}(\mu_{i};\mu_{j})=\delta_{ij}$ and
@@
Now we train a physics-reinforced neural network using DEIM, which we denote by PINN-DEIM, to solve the problem. The physics-reinforced neural network takes in the parameter for the problem and outputs an estimate of the reduced order solution vector. It is trained by minimizing the sum of the squared norm of the discretized equation residual and the squared error for the known reduced order solutions (obtained from the original POD). We similarly do this for a PINN using NEIM, which we call PINN-NEIM. Each iteration of training is slower for PINN-NEIM than for PINN-DEIM due to the need to backpropagate through the NEIM neural networks, but the loss generally decreases more per iteration for PINN-NEIM. For comparison between the two networks, however, we use 1000 training iterations and two hidden layers with 60 neurons each, for both methods.
We now compare the errors of the two methods on a test set of $m_{\text{test}}=100$ uniformly random test points. The training and test sets are shown in Figure 8.

### v2

@@ 3.4.1. Local Error Estimate
We now seek a theoretical estimate to explain the sources of error in NEIM. For simplicity, we will consider the case in which the NEIM networks approximate constant vectors, that is, $w_{t}^{(j)}(\mu_{i})>0$ if and only if $\mu_{i}=\mu^{(j)}$, and that $w_{e}(\mu_{i};\mu_{j})=\delta_{ij}$. We define the matrix of exact vectors as $M^{(k)}=\begin{bmatrix}z^{(1)}(\mu^{(1)})&\dots&z^{(k)}(\mu^{(k)})\end{bmatrix}$, and the matrix of its approximation by neural networks as $\widehat{M}^{(k)}(\mu)=\begin{bmatrix}M_{\mu^{(1)}}(\tilde{v}(\mu))&\dots&M_{\mu^{(k)}}(\tilde{v}(\mu))\end{bmatrix}$. Now define the vector of interpolated $\theta$ coefficients resulting from NEIM to be
@@
With a little abuse of notation, we denote $\operatorname{N}(\mu):=U^{\top}\operatorname{N}(v(\mu);\mu)$, and we want to estimate
@@
$\leq\|\operatorname{N}-M^{(k)}\theta^{(k)}\|_{L^{2}}+\sum_{i=1}^{k}\|z^{(i)}(\mu^{(i)})-M_{\mu^{(i)}}(\tilde{v}(\cdot))\|_{L^{2}}\|\theta_{i}^{(k)}\|_{{L^{\infty}}}$
$\hskip 4.30554pt+\sum_{i=1}^{k}\|M_{\mu^{(i)}}(\tilde{v}(\cdot))\|_{L^{2}}\|\theta_{i}^{(k)}-\widehat{\theta}_{i}^{(k)}\|_{{L^{\infty}}}$
$\leq\underbrace{\|\operatorname{N}-M^{(k)}\theta^{(k)}\|_{L^{2}}}_{\text{Projection error}}+\sum_{i=1}^{k}\underbrace{\|z^{(i)}(\mu^{(i)})-M_{\mu^{(i)}}(\tilde{v}(\cdot))\|_{L^{2}}}_{\text{Training error}}\|\theta_{i}^{(k)}\|_{{L^{\infty}}}$
$\hskip 4.30554pt+\lvert\mathcal{P}\rvert^{1/2}\sum_{i=1}^{k}\underbrace{\|\theta_{i}^{(k)}-\widehat{\theta}_{i}^{(k)}\|_{{L^{\infty}}}}_{\text{Interpolation error}}+\sum_{i=1}^{k}\underbrace{\|M_{\mu^{(i)}}(\tilde{v}(\cdot))-z^{(i)}(\mu^{(i)})\|_{L^{2}}\|\theta_{i}^{(k)}-\widehat{\theta}_{i}^{(k)}\|_{{L^{\infty}}}}_{\text{Training/interpolation error interaction}},$
where we used the triangle inequality and the fact that $\|z^{(i)}(\mu^{(i)})\|_{2}=1$.
@@ 3.5. Comparison with State-of-the-Art Techniques
Here, we qualitatively compare NEIM and the discrete empirical interpolation method (DEIM) [9] to highlight a few advantages of NEIM. We then briefly relate NEIM to DeepONets [28].
@@ 4. Numerical examples
We now present some numerical examples to demonstrate the performance of NEIM [^1] Code for all examples is available at https://github.com/maxhirsch/NEIM.. We use the open source library MLniCS [29], based on RBniCS [39] and FEniCS [27], for our physics-informed neural network example. For each experiment, we use PyTorch’s learning rate scheduler and the Adam optimizer for training the NEIM networks. The other parameters for the training and NEIM network architecture are listed in Table 2.
@@
For the liquid crystal experiment, our two NEIM approximations used the same training parameters, which we list in the “Liquid Crystal $Q$, $r$” row. The “Liquid Crystal $\theta$” row in the table lists the parameters for training the neural network which outputs the prediction of $\theta(\mu)$. We tested other interpolation methods for $\theta(\mu)$ in the other experiments depending on the complexity of the benchmark. While radial basis function interpolation could be used, it did not offer much improvement in our tests, so we used piecewise cubic and piecewise linear interpolation for the prediction of $\theta(\mu)$ instead. In what follows, the “number of modes” used in DEIM will refer to the number of components of the nonlinearity which DEIM selects to interpolate. In NEIM, the “number of modes” will simply refer to the number of terms in the NEIM expansion.
@@ 4.1. Solution Independent Function Approximation
In this first example, we approximate a parameterized function that can be thought of as the forcing term in a parameterized PDE discretization. Let us consider $s:\Omega\times\mathcal{P}\to\mathbb{R}$ defined by $s(x;\mu)=(1-x)\cos(3\pi\mu(x+1))e^{-(1+x)\mu},$ where $\Omega=[-1,1]$ and $\mu\in\mathcal{P}=[1,\pi]$, and the Poisson problem defined by
(11) $\begin{cases}\Delta v(x)=s(x;\mu),&x\in\Omega,\\ v(-1)=v(1)=0.\end{cases}$
Now let $\boldsymbol{x}=[x_{1},\dots,x_{n}]^{\top}\in\mathbb{R}^{n}$ with $x_{i}$ equally spaced points in $\Omega$ for $i=1,\dots,n$, with $n=100$, and define the forcing $f:\mathcal{P}\to\mathbb{R}^{n}$ by $f(\mu)=[s(x_{1};\mu),\dots,s(x_{n};\mu)]^{\top}\in\mathbb{R}^{n}$ for $\mu\in\mathcal{P}$. The goal is to solve the equation
@@
coming from a finite difference approximation of (11), where $v_{1}=v_{n}=0$, $h^{-2}=30$, and
@@
Figure 3. NEIM and DEIM average absolute error by number of modes in 4.1.
We now compute a NEIM approximation, where we take all weights equal to $1$, and we compare it with a DEIM approach. Because the NEIM neural networks in this case approximate a constant vector-valued function at each iteration as $f(\mu)$ is independent of $v$, we also include a comparison to “exact” NEIM, that is, the result of NEIM when the neural networks are replaced with the exact constant vectors $M_{\mu^{(i)}}(\widetilde{v})\equiv c_{i}$ for some $c_{i}\in\mathbb{R}^{r}$. This way, we can analyze the performance of NEIM in relation to the accuracy of the different components in the expansion.
In Figure 4, we see how the error in NEIM and DEIM develop as the number of modes used increases. The number of modes in DEIM is the number of singular vectors retained in the DEIM POD of the nonlinearity, while the number of modes in NEIM is the number of neural networks in the NEIM expansion. The error for all methods decreases with respect to the number of modes used. For larger numbers of modes, reaching machine precision for the singular values, DEIM outperforms NEIM, but when considering only fewer modes, NEIM performance are remarkably great over the whole parameter space. Indeed, for $r<20$ the neural empirical interpolation method shows small error at the parameters that have been selected, outperforming DEIM for many parameter values. We remark that NEIM results in small error at the selected parameters $\mu^{(j)}$ primarily because for those parameters we obtain a best fit to $v\mapsto U^{\top}\operatorname{N}(v;\mu^{(j)})$, while the nonlinearities $v\mapsto U^{\top}\operatorname{N}(v;\mu_{i})$ for $\mu_{i}\neq\mu^{(j)}$ may not be close to the span of the trained networks in the NEIM expansion. DEIM seems to have roughly the same error across the parameter space for each number of modes used, while NEIM shows error of different order of magnitude. As expected, the error for the exact version of NEIM, exploiting the exact constant vectors as modes, is smaller than that of the normal NEIM when using a large number of modes. Despite this, we remark that the accuracy of NEIM and NEIM “exact” are equivalent for $r<20$, confirming the consistency of the approach.
@@
Figure 4. NEIM and DEIM absolute error increasing number of modes in 4.1.
@@ 4.2. Solution Dependent Function Approximation
We now approximate the parameterized nonlinear function $s:\Omega\times\mathbb{R}\times\mathcal{P}\to\mathbb{R}$ defined by $s(x,v;\mu)=(1-\lvert x\rvert)e^{-(1+x)v\mu}$, where $\Omega=[-1,1]$, $\mu\in\mathcal{P}=[1,\pi]$, appearing as a nonlinear term in
(12) $\begin{cases}\Delta v(x)=s(x,v(x);\mu),\quad x\in\Omega,\\ v(-1)=v(1)=0.\end{cases}$
Now let $x_{i}$ be equally spaced points in $\Omega$ for $i=1,\dots,n$, with $n=100$, and in an abuse of notation, let $v(\mu)\in\mathbb{R}^{n}$ be the solution of the discretized system,
@@
where $h^{-2}=30$ and $f(v;\mu)$ is the nonlinear term in the system defined by $f(v;\mu)=\begin{bmatrix}s(x_{1},v_{1};\mu)&\dots&s(x_{n},v_{n};\mu)\end{bmatrix}^{\top}$. As before, we use $m=51$ equally spaced parameters in $\mathcal{P}=[1,\pi]$, and perform a POD on the matrix of solution snapshots to obtain a POD basis $U_{r}$. The goal now is to approximate $U_{r}^{\top}f(U_{r}\tilde{v};\mu)$ for $\mu\in\mathcal{P}$. Based on the plot of singular values in Figure 6, we truncate our expansion at $r=20$ modes.
@@
Figure 6. NEIM and DEIM average absolute error by number of modes in 4.2.
For this test case, we take the weights in the NEIM algorithm to be $w_{e}(\mu_{i};\mu_{j})=\delta_{ij},\quad w_{t}^{(k)}(\mu_{i})=1\quad\forall i,j,k$, and compare the accuracy results of NEIM and DEIM approaches.
@@
Now in Figure 8, we show examples of the learned modes in NEIM and DEIM. The exact nonlinearity in (a) is $U_{r}U_{r}^{\top}f(U_{r}\tilde{v}(\mu);\mu)$ for $\mu=1.72$. In (b), we see how the DEIM error is smallest at the selected interpolation points in the DEIM algorithm, as expected. Then plots (c)-(f) are the first six NEIM and DEIM modes for $\mu=1.72$ and $\mu=1.00$. In NEIM, these are each of the first six terms in the affine expansion, and in DEIM, these are the residuals approximated by each additional DEIM mode used. The NEIM and DEIM modes in (c) and (d) correspond to $\mu=1.72$, which was not one of the selected parameters in the NEIM algorithm. We see that both NEIM and DEIM have smooth approximations in the spatial domain, with primarily the first three modes in each being nonzero. Visually, we see that these modes indeed seem to add to the nonlinearity in (a). When $\mu=1$, which was the second parameter selected by the NEIM algorithm, only the first and second NEIM modes are significantly different from zero, which we expect since NEIM should almost exactly approximate the $j$th selected parameter with its first $j$ modes. It was also observed that the only significantly nonzero term in the NEIM approximation corresponding to the first selected parameter in the algorithm was the first mode.
Lastly, we compare our NEIM approximation of the nonlinearity with a simple feedforward neural network approximation of $\mu\mapsto U_{r}^{\top}f(U_{r}\tilde{v}(\mu);\mu)$ in Figure 9. We tested feedforward networks containing more parameters than NEIM and also fewer parameters than NEIM. In both cases, NEIM performed better than or comparably to the feedforward networks tested. Thus, in this example, we see that the greedy training algorithm performs better than training everything at once with a single network.
Refer to caption Refer to caption
(a) Exact Nonlinearity $\mu=1.72$ (b) NEIM/DEIM Error $\mu=1.72$
Refer to caption Refer to caption
(c) NEIM Modes $\mu=1.72$ (d) DEIM Modes $\mu=1.72$
Refer to caption Refer to caption
(e) NEIM Modes $\mu=1.00$ (f) DEIM Modes $\mu=1.00$
Figure 8. NEIM and DEIM modes in experiment 4.2.
Figure 9. Absolute error of NEIM approximation of nonlinearity versus basic neural network approximation for various architectures in experiment 4.2.
@@ 4.3. Nonlinear Elliptic Physics-Informed Neural Network
We now apply NEIM in the context of physics-informed neural networks (PINNs) for reduced order models. The goal is to combine a physics-based learning within a data-driven reduced context. More specifically, one aims at learning the parameter-to-reduced vector mapping via neural network architectures that minimize an unsupervised loss term based on the reduced residual projected onto the linear POD subspace. The setup for the problem is shown in Figure 10.
Figure 10. Setting for NEIM with physics-informed neural networks in experiment 4.3.
For further details we refer the reader to [22, 10, 33]. In order to embed the reduced physics in the loss for nonlinear problems, it is important that the approximation of the nonlinearity is compatible with automatic differentiation. When it comes to more complex nonlinearities, exploiting tensor-assembly for polynomial nonlinearities or empirical interpolation approaches can be challenging. In fact, DEIM will require assembling nonlinearities in the online phase of computation, which, if done in standard finite element packages like FEniCS [27], will not be differentiable in deep learning libraries like PyTorch [32]. On the other hand, NEIM is automatically compatible with deep learning libraries because our approach is itself a neural network with automatic differentiation included by definition.
@@
[…]
@@
Now we train a physics-informed neural network using DEIM, which we denote by PINN-DEIM, to solve the problem. The physics-informed neural network takes in the parameter for the problem and outputs an estimate of the reduced order solution vector. It is trained by minimizing the sum of the squared norm of the discretized equation residual. We similarly do this for a PINN using NEIM, which we call PINN-NEIM. Each iteration of training is slower for PINN-NEIM than for PINN-DEIM due to the need to backpropagate through the NEIM neural networks, but the loss generally decreases more per iteration for PINN-NEIM. For comparison between the two networks, however, we use 1000 training iterations and two hidden layers with 60 neurons each, for both methods.
We now compare the errors of the two methods on a test set of $m_{\text{test}}=100$ uniformly random test points. The training and test sets are shown in Figure 11. We note that in a ROM context, one would like to have the discretization of the parameter space as sparse as possible. In general, if we refine our sample for the training set, we expect that the POD would capture more information and that the greedy step in the NEIM algorithm would select more representative points with large error, leading to a more accurate approximation. In this case, we expect that the trend in the error as the number of NEIM modes increases should remain the same.